If the supply corn is Qn=a+BP and the supply by the rest of the world is Qr=c+episode what is the world supply? What is the world inverse supply?
1
Expert's answer
2015-06-09T11:21:10-0400
If the supply of corn is Qn= a + bp and the supply by the rest of the word is­ Qr=c+ep, then the world supply is Qw = Qn + Qr = a + c +(b + e)*p and the world inverse supply is p = (Qw - a - c)/(b+e).